OpenStudy (anonymous):

Find the equation of the line in slope-intercept form: 4x + 9y = 36 @TheSmartOne This one seems more tricky then the last one

TheSmartOne (thesmartone):

it's less tricky than the last one first we need to subtract 4x on both sides :D

OpenStudy (anonymous):

Okay thats what I thought was the first step. It would then be 9y = 36 + 4x or 9y = 4x + 36 then we would divide nine on both sides! Then it would be y = 4/9x + 4 right?

OpenStudy (anonymous):

Or would it be reciprical?

TheSmartOne (thesmartone):

hold on!

TheSmartOne (thesmartone):

we need to subtract 4x on both sides so we get: 9y = 36 - 4x which is the same thing as 9y = -4x + 36

OpenStudy (anonymous):

Oh so it would end of being A?

OpenStudy (anonymous):

I forgot about the negative

TheSmartOne (thesmartone):

you didn't give answer choices :P

OpenStudy (anonymous):

Oppps well answer A is y = -4/9x + 4

TheSmartOne (thesmartone):

that's correct :)